Given f(x) k(1 + x2) 2. -1 < x< 2 a) Find k. b) Calculate F(x) and the three quartiles. c) Calculate E(X), Var(X)
Answer #1

Answer: Given that f(x) = (1+st) 12222 a fiat) dx = Samlet vlak = y(x+27/5) =K*6=1 K = b. powy ( Wider-klitz) dx= x*(x+x%)*,aso. 1657 endu quartie b=1.15u 3rdu quathie c = 1.655 Ev) - I danax - Saniberse) dax=() ) = ale E (x) - J 27, el dx = 1 x 27